OpenStudy (christos):

Find the distance between the given skew lines. x=1+7t, y=3+t, z=5−3t x=4−t, y=6, z=7+2t

OpenStudy (518nad):

u=x1-x2 v=y2-y2 w=z2-z1 dist=sqrt(u^2+v^2+w^2)

OpenStudy (loser66):

Let the line one is \(L_1: \dfrac{x-1}{7}=y-3= \dfrac{z-5}{-3}\) Then, the direction of \(L_1\) is \(\vec v_1=<7,1,-3>\) and it passes through P (1,3, 5) ok?

OpenStudy (loser66):

Do the same and give me \(\vec v_2\), please

OpenStudy (loser66):

To solve this problem, you need PQ and v1, v2 cross v1, v2 to get n distance formula give you \(d =\dfrac{|PQ\bullet N|}{||N||}\)